### v1

# Decoherence-free interaction and maximally entangled state generation in double-giant-atom semi-infinite waveguide systems

Abstract
Giant atoms are artificial atoms that can couple to a waveguide non-locally. Previous work has shown that two giant atoms in a braided configuration can interact through a one-dimensional infinite waveguide, with both individual and collective atomic relaxation being fully suppressed. In this paper, however, we show that the decoherence-free interaction (DFI) between two giant atoms can be realized in both braided and nested configurations when the waveguide is semi-infinite. This protected interaction fails to appear in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. We also study the entanglement generation between two giant atoms coupled to a one-dimensional semi-infinite waveguide. The results show that the maximally entangled state is generated in both braided and nested configurations due to the formation of DFI, and in the separate configuration, the maximally achievable entanglement can exceed 0.5. This study presents a new scheme for realizing DFI and generating maximally entangled states in giant-atom waveguide-QED systems.
I INTRODUCTION
Waveguide quantum electrodynamics (QED) systems are an important platform for studying atom-photon interactions and have attracted much attention in recent years [1]. Many physical phenomena, such as single-photon transport [2, 3, 4, 5], two-photon transport [6, 7], entanglement generation between distant atoms [8, 9, 10, 11, 12, 13], and superradiant and subradiant state creation [14, 15, 16], have been observed in waveguide-QED systems. In these reports, atoms are typically idealized as point-like objects and thus they only interact with the waveguide at single points.
Recently, waveguide-QED systems containing giant atoms have garnered significant attention from many peers. In such systems, giant atoms are artificial atoms that can interact with the waveguide at multiple connection points [17]. Several interesting quantum optical phenomena have been demonstrated in giant-atom waveguide-QED systems, including frequency-dependent relaxation rates and Lamb shifts [18], non-Markovian decay dynamics [19, 20, 21], single-photon scattering [22, 23, 24], and DFI [25, 26, 27, 28, 29]. In addition, the generation [30, 31, 32, 33, 34] and transfer [35] of quantum entanglement have also been investigated in giant-atom waveguide-QED systems.
However, it is important to note that in most studies on giant atoms, the waveguide is usually considered infinite. A key question is how giant atoms behave when coupled to a semi-infinite waveguide. The semi-infinite waveguide is a one-dimensional waveguide with one end terminated by a perfect mirror [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Compared to the infinite waveguide, the semi-infinite waveguide is particularly notable in that the mirror at the end allows the atom to interact twice with the field in the waveguide. Hence, when the giant atom couples to the semi-infinite waveguide, its behavior may differ from that in the infinite waveguide. While the semi-infinite waveguide setup containing giant atoms has been addressed in previous studies [18, 49, 50], only a single giant atom was considered.
In this work, our study goes beyond the framework of a single giant atom coupled to a semi-infinite waveguide. We consider an arbitrary number of giant atoms $M$ interacting with a semi-infinite waveguide. Treating the semi-infinite waveguide as the environment of atoms, we derive the Markov quantum master equation governing the dynamical evolution of $M$ giant atoms.
We focus on the case of two giant atoms. Concretely, we analyze the behavior of two giant atoms coupled to a one-dimensional semi-infinite waveguide with three configurations: braided, separate, and nested. In doing so, we show that the DFI between two giant atoms, previously only observed within the braided configuration in infinite waveguide systems [25], can be realized in both braided and nested configurations in semi-infinite waveguide systems. This protected interaction is not supported in a semi-infinite waveguide setup involving either two separate giant atoms or two small atoms. For two small atoms coupled to a semi-infinite waveguide, it has been reported that there is no DFI between them [25].
Furthermore, we study the entanglement generation between two giant atoms coupled to a semi-infinite waveguide with three configurations. We show that the maximally achievable entanglement for all three configurations can exceed 0.5. In particular, two giant atoms in braided and nested configurations can be in a maximally entangled state, which is attributed to the formation of DFI. While the maximally entangled state has been realized in the braided configuration in infinite waveguide systems [30],
Figure 1: (Color online) Two giant atoms in a braided configuration coupled to a one-dimensional semi-infinite waveguide. The connection points are located at coordinates $x_{j1}$ and $x_{j2}$, where $j=a,b$. In our work, we consider the case where the two giant atoms share the same transition frequency, i.e., $\omega_{a}=\omega_{b}=\omega_{0}$.
the ability to generate a maximally entangled state in the nested configuration and to generate an entangled state with entanglement exceeding 0.5 in the separate configuration is unique to semi-infinite waveguide systems.
We believe that the DFI between giant atoms can find applications in various fields, e.g., in quantum simulation [51, 52, 53, 54, 55] and quantum metrology [56, 57, 58]. In these fields, protecting qubits from decoherence while enabling their interaction is crucial. Furthermore, DFI has potential applications in quantum computation [59, 60, 61, 62, 63]. For instance, future large-scale quantum computers could be operated within the DFI framework to protect the fragile quantum information during computation. It is also possible to use DFI to generate the maximally entangled state (as we discussed in this work), which is a core resource in quantum communication [64, 65] and quantum networks [66, 67].
This paper begins with Sec. II, which describes the model for giant atoms coupled to a one-dimensional semi-infinite waveguide. Then, we specifically explore the DFI and entanglement generation between two braided giant atoms in Sec. III and IV. In Sec. V, we consider the case of two giant atoms in the separate and nested configurations. Finally, we discuss the experimental feasibility of the proposed scheme and summarize the main findings of this study in Sec. VI. In addition, we provide two Appendixes: Appendix A and Appendix B. In Appendix A, the detailed derivation of the master equation is shown. In Appendix B, we present the analytical expressions for the concurrence that depend on the phase shift.
II THEORETICAL FRAMEWORK
We start by formulating a general model in which $M$ two-level giant atoms are coupled to a semi-infinite waveguide along the positive $x$ semi-axis, with each atom $j$ having $N_{j}$ connection points. In this model, we assume the semi-infinite waveguide is realized by terminating the left end (coordinate set to $x=0$) of the infinite waveguide with a mirror, and all atoms are located on the right side of the mirror.
Assuming that the coupling between giant atoms and the semi-infinite waveguide is weak at each connection point, and that the relaxation times of giant atoms are much longer than all other relevant time scales in the system, we derive the quantum master equation describing the dynamical evolution of $M$ giant atoms (see Appendix A for detailed derivations). In the interaction picture, the master equation takes the following form (hereafter, $\hbar=1$):
$\dot{\rho}=$ $-i\left[\sum_{j=1}^{M}\delta\omega_{j}\sigma_{+}^{j}\sigma_{-}^{j}+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}g_{j,k}(\sigma_{+}^{k}\sigma_{-}^{j}+\text{H.c.}),\rho\right]+\sum_{j=1}^{M}\Gamma_{j}\mathcal{D}[\sigma_{-}^{j}]\rho$ (1)
$+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}\Gamma_{coll,j,k}\left[(\sigma_{-}^{j}\rho\sigma_{+}^{k}-\frac{1}{2}\{\sigma_{+}^{k}\sigma_{-}^{j},\rho\})+\text{H.c.}\right],$
with
$\delta\omega_{j}$ $=\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{j}}\frac{\sqrt{\gamma_{jn}\gamma_{jm}}}{2}\Big{\{}\sin\left[k_{0}|x_{jn}-x_{jm}|\right]+\sin\left[k_{0}(x_{jn}+x_{jm})\right]\Big{\}},$ (2)
$g_{j,k}$ $=\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\frac{\sqrt{\gamma_{jn}\gamma_{km}}}{2}\Big{\{}\sin\left[k_{0}|x_{jn}-x_{km}|\right]+\sin\left[k_{0}(x_{jn}+x_{km})\right]\Big{\}},$
$\Gamma_{j}$ $=\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{j}}\sqrt{\gamma_{jn}\gamma_{jm}}\Big{\{}\cos\left[k_{0}|x_{jn}-x_{jm}|\right]+\cos\left[k_{0}(x_{jn}+x_{jm})\right]\Big{\}},$
$\Gamma_{coll,j,k}$ $=\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\sqrt{\gamma_{jn}\gamma_{km}}\Big{\{}\cos\left[k_{0}|x_{jn}-x_{km}|\right]+\cos\left[k_{0}(x_{jn}+x_{km})\right]\Big{\}},$
where $\delta\omega_{j}$ represents the Lamb shift of giant atom $j$, $\mathcal{D}[X]\rho=X\rho X^{\dagger}-\frac{1}{2}\{X^{\dagger}X,\rho\}$, $g_{{j,k}}$ is the exchange interaction between giant atoms, $\Gamma_{j}$ represents the individual decay rate of giant atom $j$, $\Gamma_{coll,j,k}$ is the collective decay rate for giant atoms, and H.c. denotes the Hermitian conjugate. Moreover, $\gamma_{jn}$ represents the decay rate of atom $j$ at the connection point $n$ (coordinate $x_{jn}$), $k_{0}|x_{jn}-x_{jm}|$ corresponds to the phase acquired for a photon with wave vector $k_{0}$ traveling between connection points $n$ and $m$ of the giant atom $j$, and $k_{0}(x_{jn}+x_{jm})$ represents the phase acquired during the process in which the photon released from connection point $n$ of atom $j$ is absorbed by the connection point $m$ of the same atom after being reflected through the mirror located at $x=0$, and the remaining terms can be interpreted in a similar manner.
From Eq. (2), the primary difference with respect to the infinite waveguide case [25] is that the expressions for $\delta\omega_{j}$, $g_{j,k}$, $\Gamma_{j}$, and $\Gamma_{coll,j,k}$ no longer depend just on the relative distances between the connection points: they are also determined by the distance from the connection points to the mirror in the semi-infinite waveguide case.
We focus on the case of two giant atoms where each atom interacts with the waveguide at two connection points. For two giant atoms with two connection points each, three distinct topology configurations are possible: braided, separate, and nested [25]. The setup we initially consider is the braided configuration, as shown in Fig. 1. Later, we consider the separate and nested configurations in Sec. V. To simplify the discussion, the model in Fig. 1 adopts the following assumptions: $(i)$ the decay rate at each connection point is identical and denoted by $\gamma$, and $(ii)$ the phase shift between adjacent connection points is defined as $\theta$, where $\theta/2$ represents the phase shift accumulated as a photon travels from position $x_{a1}$ to the mirror ($x=0$). When the system contains two giant atoms, Eq. (1) reduces to
$\dot{\rho}=$ $-i\left[\delta\omega_{a}\sigma_{+}^{a}\sigma_{-}^{a}+\delta\omega_{b}\sigma_{+}^{b}\sigma_{-}^{b}+g_{a,b}\left(\sigma_{+}^{b}\sigma_{-}^{a}+\text{H.c.}\right),\rho\right]$ (3)
$+\Gamma_{a}\mathcal{D}[\sigma_{-}^{a}]\rho+\Gamma_{b}\mathcal{D}[\sigma_{-}^{b}]\rho$
$+\Gamma_{coll,a,b}\left[(\sigma_{-}^{a}\rho\sigma_{+}^{b}-\frac{1}{2}\left\{\sigma_{+}^{b}\sigma_{-}^{a},\rho\right\})+\text{H.c.}\right],$
For the two braided giant atoms shown in Fig. 1, the expressions of the coefficients in Eq. (3) are given in the form below:
$\delta\omega_{a}$ $=\gamma\sin(2\theta)+\gamma\sin(3\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(5\theta),$ (4)
$\delta\omega_{b}$ $=\gamma\sin(2\theta)+\gamma\sin(5\theta)+\frac{\gamma}{2}\sin(3\theta)+\frac{\gamma}{2}\sin(7\theta),$
$g_{a,b}$ $=\frac{\gamma}{2}(3\sin\theta+\sin(3\theta)+\sin(2\theta)+2\sin(4\theta)$
$\quad+\sin(6\theta)),$
$\Gamma_{a}$ $=2\gamma+2\gamma\cos(2\theta)+2\gamma\cos(3\theta)+\gamma\cos\theta$
$\quad+\gamma\cos(5\theta),$
$\Gamma_{b}$ $=2\gamma+2\gamma\cos(2\theta)+2\gamma\cos(5\theta)+\gamma\cos(3\theta)$
$\quad+\gamma\cos(7\theta),$
$\Gamma_{coll,a,b}$ $=\gamma(3\cos\theta+\cos(3\theta)+\cos(2\theta)+2\cos(4\theta)$
$\quad+\cos(6\theta)),$
The more general forms of the coefficients, which account for arbitrary decay rates at each connection point and arbitrary phase shifts between connection points, are provided in Appendix A.
Figure 2: (Color online) Exchange interaction and decay rates as a function of $\theta/\pi$.
III DECOHERENCE-FREE INTERACTION
In previous work, it has been confirmed that two braided giant atoms can interact via the infinite waveguide without losing their excitations to it [25]. That is, there is a decoherence-free interaction (DFI) between two braided giant atoms. Mathematically, the condition for DFI to manifest can be stated as
$g_{a,b}$ $\neq 0,$ (5)
$\Gamma_{a}=\Gamma_{b}$ $=\Gamma_{coll,a,b}=0.$
[…]
From Fig. 2(a)(a), we can observe that when $\theta=\pi/2$ and $3\pi/2$, the decay rates $\Gamma_{a}$, $\Gamma_{b}$, $\Gamma_{coll,a,b}$ are all zero while the exchange interaction $g_{{a,b}}$ remains nonzero. This implies that DFI can be realized between two braided giant atoms in the semi-infinite waveguide system.
The above analysis on DFI is naturally extended to the general case of $\theta\in[n\pi,(n+1)\pi]$ (where $n$ is an integer). Based on the periodic behavior of the system, we easily find that in the region of $\theta\in[n\pi,(n+1)\pi]$, there exists DFI when
$\theta=\left(n+\frac{1}{2}\right)\pi.$ (6)
IV ENTANGLEMENT GENERATION BETWEEN TWO BRAIDED GIANT ATOMS
This section explores the entanglement generation between two braided giant atoms, where each giant atom is coupled to the semi-infinite waveguide at two distinct connection points. For two two-level atoms, the entanglement between them can be quantified by concurrence [68]. In the following, we first derive the expression for the concurrence of the two-giant-atom system.
IV.1 Concurrence
In our work, we focus on entanglement generation in the case where only one giant atom is excited. When there is only one excitation in the system, we can remove the jump term in the quantum master equation (3) to obtain a non-Hermitian effective Hamiltonian $H_{\text{eff}}$ [30, 34, 69, 70]:
$\hat{H}_{\text{eff}}$ $=\sum_{j=a,b}\delta\omega_{j}\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{j}+g_{{a,b}}(\hat{\sigma}_{+}^{b}\hat{\sigma}_{-}^{a}+\hat{\sigma}_{+}^{a}\hat{\sigma}_{-}^{b})$ (7)
$\quad-\frac{i}{2}\sum_{j=a,b}\Gamma_{j}\hat{\sigma}_{+}^{j}\hat{\sigma}_{-}^{j}-\frac{i}{2}\Gamma_{coll,a,b}(\hat{\sigma}_{+}^{b}\hat{\sigma}_{-}^{a}+\hat{\sigma}_{+}^{a}\hat{\sigma}_{-}^{b}).$
In this situation, the Schrödinger equation with the non-Hermitian Hamiltonian $H_{\text{eff}}$ can effectively describe the dynamic evolution of the two-giant-atom system
$i\frac{\partial}{\partial t}|\psi(t)\rangle=\hat{H}_{\text{eff}}|\psi(t)\rangle,$ (8)
where $|\psi(t)\rangle$ is the state vector of the system:
$|\psi(t)\rangle=c_{eg}(t)|e\rangle_{a}|g\rangle_{b}+c_{ge}(t)|g\rangle_{a}|e\rangle_{b}.$ (9)
Then, the density matrix for the two giant atoms in the basis $\left\{|e\rangle_{a}|e\rangle_{b},|e\rangle_{a}|g\rangle_{b},|g\rangle_{a}|e\rangle_{b},|g\rangle_{a}|g\rangle_{b}\right\}$ can be obtained from $\rho(t)=|\psi(t)\rangle\langle\psi(t)|$
$\rho(t)$ $=\begin{pmatrix}0&0&0&0\\ 0&|{c}_{eg}(t)|^{2}&{c}_{eg}(t){c}_{ge}^{*}(t)&0\\ 0&{c}_{eg}^{*}(t){c}_{ge}(t)&|{c}_{ge}(t)|^{2}&0\\ 0&0&0&0\end{pmatrix}$. (10)
For this density matrix, the concurrence is given by [68]:
$C(t)=2|c_{eg}(t)c_{ge}^{\ast}(t)|.$ (11)
Note that when $C=0$, the state is separable, and when $C=1$, the state is maximally entangled. In this section, and in the part of Sec. V that concerns entanglement generation, we let the two giant atoms start in the state $|e\rangle_{a}|g\rangle_{b}$. According to this initial condition, expressions for probability amplitudes $c_{\text{eg}}(t)$ and $c_{\text{ge}}(t)$ can be obtained by solving Eq. (8) and then the entanglement of two giant atoms can be measured through Eq. (11).
IV.2 Entanglement generation
We now study the entanglement generation between two braided giant atoms coupled to a semi-infinite waveguide. Based on Eqs. (7)-(11), the analytical form of concurrence is derived as follows:
$C_{B}(t)=\frac{e^{-\frac{1}{2}\,\mathrm{Re}\left[R_{1}(t)\right]}}{2}\left|\frac{R_{2}(t)\,\left(R_{3}(t)+R_{4}(t)\right)}{AA^{*}}\right|.$ (12)
where the subscript $B$ denotes the braided configuration case. Similarly, subscripts $S$ and $N$, used in Sec. V, correspond to the separate and nested configurations, respectively. The expressions for $A$ and $R_{i}(t)$ ($i=1,2,3,4$), which depend on the phase shift $\theta$, are given in Appendix B.
[…]
To clearly elucidate the phase-shift dependence of entanglement generation, we now analyze the evolution of the concurrence $C_{B}$ when phase shift $\theta$ takes some specific values, where $\theta\in[0,\pi]$. As shown in Fig. 3(a)(a), when $\theta=0$, the concurrence $C_{B}$ exhibits steady-state behaviors.
Figure 3: (Color online) $(a)$ Concurrence $C_{B}$ as a function of $\theta/\pi$ and $\gamma t$. $(b)$ Concurrence $C_{B}(t)$ as a function of $\gamma t$ at a given value of phase shift.
The steady-state behavior of the concurrence $C_{B}$ at $\theta=0$ can be explained through the analytical expression. When $\theta=0$, the concurrence in Eq. (12) becomes
$C_{B}(t)=\frac{1}{2}\left({1-e^{-16\gamma t}}\right).$ (13)
[…]
$C_{B}(t)=|\sin(2\gamma t)|.$ (14)
According to Eq. (14), it can be found that the concurrence $C_{B}(t)$ oscillates between zero and one with a period $\frac{\pi}{2\gamma}$. When $t=\frac{n\pi}{4\gamma}$ (where $n$ is an odd integer), concurrence $C_{B}(t)=1$, indicating that the two giant atoms reach a maximally entangled state. Specifically, when $\theta$ is taken to be $\pi$, no entanglement is generated here, i.e., $C_{B}=0$. The reason is mainly that when $\theta=\pi$, exchange interaction $g_{a,b}$, individual decay rates $\Gamma_{a}$ and $\Gamma_{b}$, as well as the collective decay rate $\Gamma_{\text{coll},a,b}$ all become zero (as can be seen from Fig. 2), which implies that the two giant atoms decouple from the waveguide, thereby making entanglement generation impossible.
Figure 4: (Color online) $(a)$ Two giant atoms in a separate configuration coupled to a one-dimensional semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$. $(c)$ Concurrence $C_{s}$ as a function of $\theta/\pi$ and $\gamma t$.
Figure 5: (Color online) $(a)$ Two giant atoms in a nested configuration coupled to a one-dimensional semi-infinite waveguide. $(b)$ Exchange interaction and decay rates as a function of $\theta/\pi$. $(c)$ Concurrence $C_{N}$ as a function of $\theta/\pi$ and $\gamma t$.
V TWO GIANT ATOMS IN SEPARATE AND NESTED CONFIGURATIONS
In the above sections, we have analyzed the behavior of two braided giant atoms coupled to a semi-infinite waveguide, as well as their entanglement properties. In this section, as shown in Fig. 4 and Fig. 5, we consider the separate and nested configurations. According to Eq. (2), the relevant parameters for separate [Fig. 4(a)(a)] and nested [Fig. 5(a)(a)] configurations can be obtained as
$\delta\omega_{a}$ $=\gamma\sin\theta+\gamma\sin(2\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(3\theta),$ (15)
$\delta\omega_{b}$ $=\gamma\sin\theta+\gamma\sin(6\theta)+\frac{\gamma}{2}\sin(5\theta)+\frac{\gamma}{2}\sin(7\theta),$
$g_{a,b}$ $=\frac{\gamma}{2}(\sin\theta+2\sin(2\theta)+2\sin(3\theta)+2\sin(4\theta)$
$\quad+\sin(5\theta)),$
$\Gamma_{a}$ $=2\gamma+2\gamma\cos\theta+\gamma\cos\theta+\gamma\cos(3\theta)$
$\quad+2\gamma\cos(2\theta),$
$\Gamma_{b}$ $=2\gamma+2\gamma\cos\theta+\gamma\cos(5\theta)+\gamma\cos(7\theta)$
$\quad+2\gamma\cos(6\theta),$
$\Gamma_{coll,a,b}$ $=\gamma(\cos\theta+2\cos(2\theta)+2\cos(3\theta)+2\cos(4\theta)$
$\quad+\cos(5\theta)).$
$\delta\omega_{a}$ $=\gamma\sin(3\theta)+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin\theta+\frac{\gamma}{2}\sin(7\theta),$ (16)
$\delta\omega_{b}$ $=\gamma\sin\theta+\gamma\sin(4\theta)+\frac{\gamma}{2}\sin(3\theta)+\frac{\gamma}{2}\sin(5\theta),$
$g_{a,b}$ $=\frac{\gamma}{2}(2\sin\theta+3\sin(2\theta)+\sin(3\theta)+\sin(5\theta)$
$\quad+\sin(6\theta)),$
$\Gamma_{a}$ $=2\gamma+2\gamma\cos(3\theta)+2\gamma\cos(4\theta)+\gamma\cos\theta$
$\quad+\gamma\cos(7\theta),$
$\Gamma_{b}$ $=2\gamma+2\gamma\cos\theta+2\gamma\cos(4\theta)+\gamma\cos(3\theta)$
$\quad+\gamma\cos(5\theta),$
$\Gamma_{coll,a,b}$ $=\gamma(2\cos\theta+3\cos(2\theta)+\cos(3\theta)+\cos(5\theta)$
$\quad+\cos(6\theta)).$
Using a strategy similar to that used in the braided configuration, we conclude that the dependence of exchange interaction and decay rates on $\theta$ in Eqs. (15) and (16) is a $2\pi$-periodic function. Therefore, we restrict our discussion to the region $\theta\in[0,2\pi]$, which fully captures the behavior of the system. In Fig. 4(b)(b) and Fig. 5(b)(b) we show the exchange interaction and decay rates from Eqs. (15) and (16) as a function of $\theta/\pi$, where $\theta\in[0,2\pi]$.
For the separate configuration, we observe from Fig. 4(b)(b) that there is no value of $\theta$ that satisfies the DFI condition (5). However, for the nested configuration, we note that the DFI condition (5) is satisfied when $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$, as shown in Fig. 5(b)(b). The result demonstrates that our scheme enables DFI in the nested configuration, challenging the prevailing notion that DFI can only be achieved in the braided configuration [25, 26, 27, 28, 29].
Similar to the braided configuration case, we generalize the study of DFI in the nested configuration to the region of $\theta\in[n\pi,(n+1)\pi]$, where $n$ is an integer. Given the periodicity of the system, the phase shift for the manifestation of DFI in the nested configuration in the region of $\theta\in[n\pi,(n+1)\pi]$ can be easily determined and is given by:
$\theta=\left(n+\frac{2k+1}{4}\right)\pi,\quad k=0,1.$ (17)
The concurrences $C_{s}$ and $C_{N}$ for two giant atoms in separate and nested configurations as a function of $\theta/\pi$ and $\gamma t$ are plotted in Fig. 4(c)(c) and Fig. 5(c)(c), respectively. From Fig. 4(c)(c) and Fig. 5(c)(c), we observe that both concurrences $C_{s}$ and $C_{N}$ reach a steady-state value at $\theta=0$ and $2\pi$, but drop to zero at $\theta=\pi$. These behaviors are consistent with those observed in the infinite waveguide scheme [30].
In particular, the value of concurrence $C_{s}$ can exceed 0.5 when $\theta\rightarrow\pi/2$ and $3\pi/2$, and the concurrence $C_{N}$ is characterized by periodic oscillations ranging from zero to one at the values of $\theta$ where DFI manifests (i.e., $\theta=\pi/4,3\pi/4,5\pi/4,7\pi/4$). This indicates that two nested giant atoms in the semi-infinite waveguide system can also be prepared in the maximally entangled state. Notably, in the infinite waveguide scheme [30], it is not possible to generate entangled states with concurrence exceeding 0.5 in the separate configuration as well as maximally entangled states in the nested configuration.
VI DISCUSSION AND CONCLUSION
We present a discussion on the experimental feasibility of the proposed scheme. The giant atom setup has been experimentally realized recently [29, 71, 72, 73]. In particular, Ref. [29] reports a realization that two giant atoms, whose connection points are arranged in a braided manner, are coupled to an infinite microwave waveguide. Furthermore, the coupling of two small atoms to a microwave waveguide with one end terminated by a mirror has been demonstrated in experiment [74]. These results suggest that our scheme is achievable with current and near-future experimental conditions.
To summarize, we have derived the quantum master equation for multiple giant atoms coupled to a semi-infinite waveguide, where each atom has multiple connection points. We specifically analyzed the behavior of two giant atoms coupled to a semi-infinite waveguide. Concretely, we considered three configurations: separate, braided, and nested. We show that two giant atoms in braided and nested configurations can interact through the semi-infinite waveguide without decohering, i.e., there is a DFI between two giant atoms. Such DFI is not possible in semi-infinite waveguide systems containing two separate giant atoms or two small atoms. This finding goes beyond the result shown in Ref. [25] for the case of an infinite waveguide, where DFI can only be realized in the braided configuration.
Moreover, we study the generation of entanglement between two giant atoms under three configurations in the semi-infinite waveguide system. In braided and nested configurations, the two giant atoms can reach a maximally entangled state due to the formation of DFI. In the separate configuration, the generated entanglement between the two giant atoms can exceed 0.5. Compared to the infinite waveguide scheme [30], the maximally achievable entanglement for two giant atoms in separate and nested configurations is enhanced in the semi-infinite waveguide system.
This work not only verifies the possibility of realizing DFI in a semi-infinite waveguide but also provides a theoretical foundation for quantum information processing based on giant-atom semi-infinite waveguide-QED systems.
ACKNOWLEDGMENTS
This work was supported by National Natural Science Foundation of China(Grants No. 11874190 and No. 12247101). Support was also provided by Supercomputing Center of Lanzhou University.
Appendix
Appendix A Derivation of the quantum master equation for multiple giant atoms with multiple connection points
A.1 Two braided giant atoms with two connection points each
Figure 6: (Color online) $(a)$ A sketch that shows the parameters of two braided giant atoms in a semi-infinite waveguide. $(b)$ Diagram illustrating the SLH calculation for two braided giant atoms in a semi-infinite waveguide.
We derived the quantum master equation for two braided giant atoms, each coupled to a one-dimensional semi-infinite waveguide at two connection points, using the $(S,L,H)$ formalism [75, 76, 77]. In this formalism, $n$ input-output ports are needed to find the triplet for a system: S (an $n\times n$ scattering matrix), $L$ (a vector consisting of $n$ matrices that describe the system-environment interaction), and the Hamiltonian $H$ of the system (including atom-atom interaction).
Once the SLH triplet G = ($\mathbf{S}$, $\mathbf{L}$, H) for a system is calculated, the master equation for the system can be easily obtained,
$\dot{\rho}=-i[H,\rho]+\sum_{k=1}^{n}\mathcal{D}[L_{k}]\rho,$ (18)
where $\mathcal{D}[L]\rho=L\rho L^{\dagger}-\frac{1}{2}\left(L^{\dagger}L\rho+\rho L^{\dagger}L\right)$ are Lindblad operator.
To achieve this, we present the two braided ginat atom setup and corresponding schematic representation of the SLH calculation in Fig. 6. The triplets for the right moving part and the left moving part are respectively
$G_{\text{R}}$ $=G_{b\text{2R}}\triangleleft G_{\theta_{3}}\triangleleft G_{a\text{2R}}\triangleleft G_{\theta_{2}}\triangleleft G_{b\text{1R}}\triangleleft G_{\theta_{1}}\triangleleft G_{a\text{1R}},$
$G_{\text{L}}$ $=G_{a\text{1L}}\triangleleft G_{\theta_{1}}\triangleleft G_{b\text{1L}}\triangleleft G_{\theta_{2}}\triangleleft G_{a\text{2L}}\triangleleft G_{\theta_{3}}\triangleleft G_{b\text{2L}},$ (19)
where
$G_{a1R}$ $=\left(1,\sqrt{\frac{\gamma_{1}}{2}}\sigma_{-}^{a},\omega_{a}\frac{\sigma_{z}^{a}}{2}\right),$ $\quad G_{a2R}$ $=\left(1,\sqrt{\frac{\gamma_{3}}{2}}\sigma_{-}^{a},0\right),$ (20)
$G_{a1L}$ $=\left(1,\sqrt{\frac{\gamma_{1}}{2}}\sigma_{-}^{a},0\right),$ $\quad G_{a2L}$ $=\left(1,\sqrt{\frac{\gamma_{3}}{2}}\sigma_{-}^{a},0\right),$
$G_{b1R}$ $=\left(1,\sqrt{\frac{\gamma_{2}}{2}}\sigma_{-}^{b},\omega_{b}\frac{\sigma_{z}^{b}}{2}\right),$ $\quad G_{b2R}$ $=\left(1,\sqrt{\frac{\gamma_{4}}{2}}\sigma_{-}^{b},0\right),$
$G_{b1L}$ $=\left(1,\sqrt{\frac{\gamma_{2}}{2}}\sigma_{-}^{b},0\right),$ $\quad G_{b2L}$ $=\left(1,\sqrt{\frac{\gamma_{4}}{2}}\sigma_{-}^{b},0\right).$
and the phase shift acquired is represented by $G_{\theta_{k}}=(e^{i\theta_{k}},0,0)$, where $k=1,2,3$ (Specifically, $G_{\theta}=(e^{i\theta},0,0)$). The connection between two SLH triplets, symbolized as $G_{2}\triangleleft G_{1}$, results in the following outcome [75, 76, 77]:
$G_{2}\triangleleft G_{1}$ $=(\mathbf{S_{2}},\mathbf{L_{2}},H_{2})\triangleleft(\mathbf{S_{1}},\mathbf{L_{1}},H_{1})$ (21)
$=\left(\mathbf{S}_{2}\mathbf{S}_{1},\mathbf{S}_{2}\mathbf{L}_{1}+\mathbf{L}_{2},H_{1}+H_{2}+\frac{1}{2i}\left[\mathbf{L}_{2}^{\dagger}\mathbf{S}_{2}\mathbf{L}_{1}-\mathbf{L}_{1}^{\dagger}\mathbf{S}_{2}^{\dagger}\mathbf{L}_{2}\right]\right)$
Additionally, each SLH triplet in Eq. (20) is given according to the definition of SLH triplet at each connection point [18]:
[…]
$G_{L}$ $=\begin{cases}(1,\sqrt{\frac{\gamma_{jn}}{2}}\sigma_{-}^{j},0)&\hskip 34.14322pt\text{applicable to all connection points of atom }j.\end{cases}$ (A.4)
Then, the total triplet for the system is
$G_{\text{tot}}$ $=(\mathbf{S_{tot}},\mathbf{L_{tot}},H_{tot})=G_{\text{R}}\triangleleft G_{\theta}\triangleleft G_{\text{L}}$ (22)
with
$S_{\text{tot}}$ $=e^{i(2\theta_{1}+2\theta_{2}+2\theta_{3}+\theta)},$ (23)
$L_{\text{tot}}$ $=\left((e^{i(\theta_{1}+\theta_{2}+\theta_{3}+\theta)}+e^{i(\theta_{1}+\theta_{2}+\theta_{3})})\sqrt{\frac{\gamma_{1}}{2}}+(e^{i(2\theta_{1}+2\theta_{2}+\theta_{3}+\theta)}+e^{i\theta_{3}})\sqrt{\frac{\gamma_{3}}{2}}\right)\sigma_{-}^{a}$
$+\left((e^{i(2\theta_{1}+\theta_{2}+\theta_{3}+\theta)}+e^{i(\theta_{2}+\theta_{3})})\sqrt{\frac{\gamma_{2}}{2}}+(e^{i(2\theta_{1}+2\theta_{2}+2\theta_{3}+\theta)}+1)\sqrt{\frac{\gamma_{4}}{2}}\right)\sigma_{-}^{b},$
$H_{\text{tot}}$ $=\omega_{a}\frac{\sigma_{z}^{a}}{2}+\omega_{b}\frac{\sigma_{z}^{b}}{2}+\left[\sqrt{\gamma_{1}\gamma_{3}}(\sin(\theta_{1}+\theta_{2})+\sin(\theta_{1}+\theta_{2}+\theta))+\frac{\gamma_{1}}{2}\sin\theta+\frac{\gamma_{3}}{2}\sin(2\theta_{1}+2\theta_{2}+\theta)\right]\sigma_{+}^{a}\sigma_{-}^{a}$
$+\left[\sqrt{\gamma_{2}\gamma_{4}}(\sin(\theta_{2}+\theta_{3})+\sin(2\theta_{1}+\theta_{2}+\theta_{3}+\theta))+\frac{\gamma_{2}}{2}\sin(2\theta_{1}+\theta)+\frac{\gamma_{4}}{2}\sin(2\theta_{1}+2\theta_{2}+2\theta_{3}+\theta)\right]\sigma_{+}^{b}\sigma_{-}^{b}$
$+\left[\frac{\sqrt{\gamma_{1}\gamma_{2}}}{2}(\sin\theta_{1}+\sin(\theta_{1}+\theta))+\frac{\sqrt{\gamma_{2}\gamma_{3}}}{2}(\sin\theta_{2}+\sin(2\theta_{1}+\theta_{2}+\theta))\right$.
$\left.+\frac{\sqrt{\gamma_{3}\gamma_{4}}}{2}(\sin\theta_{3}+\sin(2\theta_{1}+2\theta_{2}+\theta_{3}+\theta))+\frac{\sqrt{\gamma_{1}\gamma_{4}}}{2}(\sin(\theta_{1}+\theta_{2}+\theta_{3})+\sin(\theta_{1}+\theta_{2}+\theta_{3}+\theta))\right]$
$\quad\left(\sigma_{+}^{b}\sigma_{-}^{a}+\sigma_{+}^{a}\sigma_{-}^{b}\right).$
Using the following properties of the $\bm{L}$ operator:
$D[a+b]\rho$ $=(a+b)\rho(a^{\dagger}+b^{\dagger})-\frac{1}{2}\left[(a^{\dagger}+b^{\dagger})(a+b)\rho+\rho(a^{\dagger}+b^{\dagger})(a+b)\right]$
$=a\rho a^{\dagger}+a\rho b^{\dagger}+b\rho a^{\dagger}+b\rho b^{\dagger}-\frac{1}{2}\left[(a^{\dagger}a+a^{\dagger}b+b^{\dagger}a+b^{\dagger}b)\rho+\rho(a^{\dagger}a+a^{\dagger}b+b^{\dagger}a+b^{\dagger}b)\right]$
$=D[a]\rho+D[b]\rho+a\rho b^{\dagger}+b\rho a^{\dagger}-\frac{1}{2}\left(a^{\dagger}b+b^{\dagger}a\right)\rho-\frac{1}{2}\rho\left(a^{\dagger}b+b^{\dagger}a\right).$ (A.6)
The master equation finally reads as
$\dot{\rho}$ $=-i[H_{\text{tot}},\rho]+\mathcal{D}\left[L_{\text{tot}}\right]\rho$
$=-i[H_{\text{tot}},\rho]+\mathcal{D}\left[\left((e^{i(\theta_{1}+\theta_{2}+\theta_{3}+\theta)}+e^{i(\theta_{1}+\theta_{2}+\theta_{3})})\sqrt{\frac{\gamma_{1}}{2}}+(e^{i(2\theta_{1}+2\theta_{2}+\theta_{3}+\theta)}+e^{i\theta_{3}})\sqrt{\frac{\gamma_{3}}{2}}\right)\sigma_{-}^{a}\right$.
$\quad+\left.\left((e^{i(2\theta_{1}+\theta_{2}+\theta_{3}+\theta)}+e^{i(\theta_{2}+\theta_{3})})\sqrt{\frac{\gamma_{2}}{2}}+(e^{i(2\theta_{1}+2\theta_{2}+2\theta_{3}+\theta)}+1)\sqrt{\frac{\gamma_{4}}{2}}\right)\sigma_{-}^{b}\right]\rho$
$=-i[H_{\text{tot}},\rho]+\left[\gamma_{1}+\gamma_{3}+\gamma_{1}\cos\theta+\gamma_{3}\cos(2\theta_{1}+2\theta_{2}+\theta)+2\sqrt{\gamma_{1}\gamma_{3}}\cos(\theta_{1}+\theta_{2})+2\sqrt{\gamma_{1}\gamma_{3}}\cos(\theta_{1}+\theta_{2}+\theta)\right]\mathcal{D}[\sigma_{-}^{a}]\rho$
$+\left[\gamma_{2}+\gamma_{4}+\gamma_{2}\cos(2\theta_{1}+\theta)+\gamma_{4}\cos(2\theta_{1}+2\theta_{2}+2\theta_{3}+\theta)+2\sqrt{\gamma_{2}\gamma_{4}}\cos(\theta_{2}+\theta_{3})+2\sqrt{\gamma_{2}\gamma_{4}}\cos(2\theta_{1}+\theta_{2}+\theta_{3}+\theta)\right]\mathcal{D}[\sigma_{-}^{b}]\rho$
$+\left[\sqrt{\gamma_{1}\gamma_{2}}(\cos\theta_{1}+\cos(\theta_{1}+\theta))+\sqrt{\gamma_{2}\gamma_{3}}(\cos\theta_{2}+\cos(2\theta_{1}+\theta_{2}+\theta))\right$.
$\left.+\sqrt{\gamma_{3}\gamma_{4}}(\cos\theta_{3}+\cos(2\theta_{1}+2\theta_{2}+\theta_{3}+\theta))+\sqrt{\gamma_{1}\gamma_{4}}(\cos(\theta_{1}+\theta_{2}+\theta_{3})+\cos(\theta_{1}+\theta_{2}+\theta_{3}+\theta))\right]$
$\quad\times\left[(\sigma_{-}^{a}\rho\sigma_{+}^{b}-\frac{1}{2}\left\{\sigma_{+}^{b}\sigma_{-}^{a},\rho\right\})+\text{H.c.}\right].$ (A.7)
In particular, when Eq. (A.7) is transformed into the interaction picture and the assumptions $\gamma_{1}=\gamma_{2}=\gamma_{3}=\gamma_{4}=\gamma$, $\theta_{1}=\theta_{2}=\theta_{3}=\theta$ are made, Eq. (A.7) reduces to Eq. (3) of the main text, with the coefficients given in Eq. (4).
A.2 Multiple giant atoms with multiple connection points each
The master equation for $M$ giant atoms coupled to a semi-infinite waveguide, where each atom $j$ has $N_{j}$ connection points, can be derived by following the same procedure described above. The total SLH triplet for the system is given by
$S_{\text{tot}}$ $=e^{ik_{0}\left[|x_{1_{1}}-x_{M_{N}}|+(x_{1_{1}}+x_{M_{N}})\right]},$ (24)
$L_{\text{tot}}$ $=\sum_{j=1}^{M}\sum_{n=1}^{N_{j}}\Big{\{}e^{ik_{0}|x_{jn}-x_{M_{N}}|}+e^{ik_{0}(x_{jn}+x_{M_{N}})}\Big{\}}\sqrt{\frac{\gamma_{jn}}{2}}\sigma_{-}^{j},$
$H_{\text{tot}}$ $=\sum_{j=1}^{M}\omega_{j}\frac{\sigma_{z}^{j}}{2}+\sum_{j=1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{j}}\frac{\sqrt{\gamma_{jn}\gamma_{jm}}}{2}\Big{\{}\sin\left[k_{0}|x_{jn}-x_{jm}|\right]+\sin\left[k_{0}(x_{jn}+x_{jm})\right]\Big{\}}\sigma_{+}^{j}\sigma_{-}^{j}$
$+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\frac{\sqrt{\gamma_{jn}\gamma_{km}}}{2}\Big{\{}\sin\left[k_{0}|x_{jn}-x_{km}|\right]+\sin\left[k_{0}(x_{jn}+x_{km})\right]\Big{\}}\left(\sigma_{+}^{(k)}\sigma_{-}^{(j)}+\text{H.c.}\right).$
where $x_{1_{1}}$ and $x_{M_{N}}$ denote the coordinates of the connection points that are closest to and farthest from the mirror respectively.
From the SLH triplet above, the Lindblad master equation including $M$ giant atoms can be formulated as
$\dot{\rho}$ $=-i[H_{\text{tot}},\rho]+\mathcal{D}\left[L_{tot}\right]\rho$
$=-i[H_{\text{tot}},\rho]+\mathcal{D}\left[\sum_{j=1}^{M}\sum_{n=1}^{N_{j}}\Big{\{}e^{ik_{0}|x_{jn}-x_{M_{N}}|}+e^{ik_{0}(x_{jn}+x_{M_{N}})}\Big{\}}\sqrt{\frac{\gamma_{jn}}{2}}\sigma_{-}^{j}\right]\rho.$ (A9)
By expanding the second term on the right-hand side of Eq. (A9) with the properties of the $\bm{L}$ operator
$D\left[\sum_{j=1}^{M}O_{j}\right]\rho=\sum_{j=1}^{M}\sum_{i=1}^{M}O_{j}\rho O_{i}^{\dagger}-\frac{1}{2}\sum_{j=1}^{M}\sum_{i=1}^{M}\left(O_{i}^{\dagger}O_{j}\rho+\rho O_{i}^{\dagger}O_{j}\right),$ (25)
we finally arrive at
$\dot{\rho}$ $=-i[H_{\text{tot}},\rho]+\mathcal{D}\left[\sum_{j=1}^{M}\sum_{n=1}^{N_{j}}\Big{\{}e^{ik_{0}|x_{jn}-x_{M_{N}}|}+e^{ik_{0}(x_{jn}+x_{M_{N}})}\Big{\}}\sqrt{\frac{\gamma_{jn}}{2}}\sigma_{-}^{j}\right]\rho$
$=-i[H_{\text{tot}},\rho]+\sum_{j=1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{j}}\sqrt{\gamma_{j_{n}}\gamma_{j_{m}}}\Big{\{}\cos\left[k_{0}|x_{jn}-x_{jm}|\right]+\cos\left[k_{0}(x_{jn}+x_{jm})\right]\Big{\}}\mathcal{D}[\sigma_{-}^{j}]\rho$
$+\sum_{j=1}^{M-1}\sum_{k=j+1}^{M}\sum_{n=1}^{N_{j}}\sum_{m=1}^{N_{k}}\sqrt{\gamma_{j_{n}}\gamma_{k_{m}}}\Big{\{}\cos\left[k_{0}|x_{jn}-x_{km}|\right]+\cos\left[k_{0}(x_{jn}+x_{km})\right]\Big{\}}\left[\left(\sigma_{-}^{(j)}\rho\sigma_{+}^{(k)}-\frac{1}{2}\left\{\sigma_{+}^{(k)}\sigma_{-}^{(j)},\rho\right\}\right)+\text{H.c.}\right].$ (A11)
When we transform Eq. (A11) into the interaction picture, it reduces to Eq. (1) in the main text, with the corresponding coefficients provided in Eq. (2).
Appendix B Analytical forms of the Phase-Dependent functions $R_{i}(t)$ $(i=1,2,3,4)$ in Eq. (12)
Expressions of the functions $R_{i}(t)$ $(i=1,2,3,4)$ in Eq. (12) have the following form:
$R_{1}(t)$ $=\gamma t\left(D+2A\cos(\frac{\theta}{2})\right),$ (26)
$R_{2}(t)$ $=E\left(e^{\sqrt{2}\,B^{*}\gamma t\cos\left(\frac{\theta}{2}\right)}-1\right),$
$R_{3}(t)$ $=2A\left(e^{\sqrt{2}\,B\gamma t\cos\left(\frac{\theta}{2}\right)}+1\right),$
$R_{4}(t)$ $=F\left(e^{\sqrt{2}\,B\gamma t\cos\left(\frac{\theta}{2}\right)}-1\right),$
with
$A$ $=\sqrt{e^{3i\theta}\left(37-50e^{i\theta}+93e^{2i\theta}-80e^{3i\theta}+86e^{4i\theta}-52e^{5i\theta}+38e^{6i\theta}-16e^{7i\theta}+9e^{8i\theta}-2e^{9i\theta}+e^{10i\theta}\right)},$ (27)
$D$ $=4+e^{i\theta}+4e^{2i\theta}+3e^{3i\theta}+3e^{5i\theta}+e^{7i\theta},$
$E$ $=-3+\cos{\theta}\left(3-4i\sin{\theta}\right)+i\sin{\theta}-4\cos{2\theta},$
[…]
$B$ $=\sqrt{(B_{1}+B_{2})e^{8i\theta}}.$
Here, $B_{1}$ and $B_{2}$ are defined as
[…]